# Utilizing Large Language Models to Optimize the Detection and Explainability of Phishing Websites

### v1

@@ 7.2. Choice of generative model
For automated explainable blocklisting, it is crucial that the AI model can generate coherent and contextually appropriate human dialogue. Our decision to use GPT 3.5T, a powerful open-domain chatbot, was informed by several limitations observed in locally deployable LLMs. Firstly, while the DistilBERT architecture is adequate for binary classification tasks, its capabilities in text generation (along with its parent model - BERT) are insufficient (Chan and Fan, 2019). More robust models such as LLaMA 2 and T5, although capable of producing meaningful explanations, were deemed impractical for real-world application due to their high resource consumption and slow performance. For instance, the smallest LLaMA 2 model, with 7 billion parameters (TheBloke, [n. d.]), required an average of 72 seconds to generate an explanation on our test system. In contrast, GPT-3.5T not only outperforms these local LLMs in both speed and output quality—taking an average of just 4 seconds to generate an explanation—but it also presents a more cost-effective solution. The operational cost of GPT-3.5T is significantly lower than that of the GPT-4 model, averaging about \$1 for 2,000 explanations.
@@ 7.3. Reliability of generated content
Pre-processing content: To ensure consistency in GPT-3.5T’s output, we prompted it to compile responses as a structured JSON with key-value pairs for the feature name, location, and description (i.e., the context of the suspicious feature within the website). This data is parsed by a driver to generate a warning displayed to the user. We utilized our codebook (https://tinyurl.com/vcjer5my) to define specific feature terms, instructing GPT-3.5T to select from this predefined list, ensuring the model does not invent alternate terms for the same feature. Only visible website features are considered, omitting features like ”Session Hijacking attempt,” ”IP Tracking,” or ”JavaScript redirection.” If GPT-3.5T identifies features not in our list but significantly supporting detection, we encouraged their identification. During the evaluation, GPT-3.5T identified three new features: ”Grammatical errors” (mistakes in spelling, grammar, and punctuation), ”Poor security indicators” (e.g., use of HTTP instead of HTTPS), and ”Mismatched URLs” (URLs not matching the official URL of the targeted organization). These were subsequently added to our feature list.
@@
Human review: LLMs can hallucinate information. Despite our structured prompt, we needed to evaluate the reliability of ChatGPT’s output. Two coders manually reviewed a sample of explainable warnings generated for 100 randomly selected websites from our training dataset. Coders were provided with a screenshot of the website, its parsed source code, and its URL. They assessed: a) if GPT-3.5T’s features were present on the website, b) if the feature location was correct, c) if the feature description was accurate, and d) if any additional features were missed. The coders achieved a Cohen’s Kappa inter-reliability agreement of 0.78. After reviewing the websites, they resolved disagreements through discussion. In total, 401 feature points were identified across 100 websites. The targeted organization imitated by the phishing website was correctly identified in all cases. Coders found 383 ( 96%) features correctly identified, with the correct location for 339 ( 84%) of them. Among the 18 incorrect identifications were non-sensitive text fields (9), hallucinated two-factor authentication modules (5), and multiple brand texts (4).
8. Discussions and Future Work
Applications
We open-source our framework (https://tinyurl.com/vcjer5my and also release an extension that supports Chromium-based web browsers which is light on resources, and protects users in real time against visiting phishing websites. While the extension does not maintain a blocklist, it does have a cache to reduce predictions on frequently opened websites, as well as a frequently updated whitelist from Tranco (Pochat et al., 2018). We also do not retain any browser history from the user except websites that are detected as malicious; to ensure that a previously detected website is not reevaluated on our server. We provide a website (https://freephish.org) where users scan any URL using the PhishLang model and also provide a list of URLs that PhishLang has detected over time. Figure 10 illustrates the main page of the site. Along with the malicious URL, the blocklist provides its IP address, the time when it was discovered, the confidence score of our model for that URL, a screenshot of the website. Thus the blocklist provides an additional resource that can be used by organizations and researchers alike to obtain phishing metadata. Figure 10 shows the Online blocklist in action.
Impact of explainable blocklisting
We plan to follow our findings in Section 7 by conducting a comprehensive user study to identify how these warnings influence user trust and promote phishing awareness. This will involve investigating user interactions with the warnings, understanding their responses, and measuring the short-term and long-term impact on their decision-making processes in various scenarios.
Figure 10. Example of Phishing blocking a phishing website. Note the website URLs have not been redacted since they are already inactive.
@@ 9. Conclusion
We introduced PhishLang, a novel phishing detection framework leveraging a language model to identify phishing websites by comparing source codes with established phishing patterns. PhishLang outperforms various machine learning and deep learning models, reducing detection time and resources while showing strong resilience against adversarial attacks. In real-time, it identified several new phishing websites, including evasive attacks, missed by popular anti-phishing tools. Reporting these threats improved detection rates, demonstrating PhishLang’s impact. We integrated PhishLang with GPT-3.5 Turbo for ”Explainable Blocklisting,” providing contextual explanations for phishing alerts, enhancing user engagement and understanding. PhishLang is open-sourced and available to users via a Chromium-based browser extension and a URL-scanning website.

### v2

@@ 6.2. Choice of generative model
For automated explainable blocklisting, it is crucial that the AI model can generate coherent and contextually appropriate human dialogue. Our decision to use GPT 3.5T, a powerful open-domain chatbot, was informed by several limitations observed in locally deployable LLMs. Firstly, while the MobileBERT architecture is adequate for binary classification tasks, its capabilities in text generation (along with its parent model - BERT) are insufficient (Chan and Fan, 2019). More robust models such as LLaMA 2 and T5, although capable of producing meaningful explanations, were deemed impractical for real-world application due to their high resource consumption and slow performance. For instance, the smallest LLaMA 2 model, with 7 billion parameters (TheBloke, [n. d.]), required an average of 72 seconds to generate an explanation on our test system. In contrast, GPT-3.5T not only outperforms these local LLMs in both speed and output quality—taking an average of just 4 seconds to generate an explanation—but it also presents a more cost-effective solution. The operational cost of GPT-3.5T is significantly lower than that of the GPT-4 model, averaging about \$1 for 2,000 explanations.
@@ 6.3. Reliability of generated content
Pre-processing content: To ensure consistency in GPT-3.5T’s output, we prompted it to compile responses as a structured JSON with key-value pairs for the feature name, location, and description (i.e., the context of the suspicious feature within the website). This data is parsed by a driver to generate a warning displayed to the user. We utilized our codebook (https://github.com/UTA-SPRLab/phishlang) to define specific feature terms, instructing GPT-3.5T to select from this predefined list, ensuring the model does not invent alternate terms for the same feature. Only visible website features are considered, omitting features like ”Session Hijacking attempt,” ”IP Tracking,” or ”JavaScript redirection.” If GPT-3.5T identifies features not in our list but significantly supporting detection, we encouraged their identification. During the evaluation, GPT-3.5T identified three new features: ”Grammatical errors” (mistakes in spelling, grammar, and punctuation), ”Poor security indicators” (e.g., use of HTTP instead of HTTPS), and ”Mismatched URLs” (URLs not matching the official URL of the targeted organization). These were subsequently added to our feature list.
@@
Human review: LLMs can hallucinate information (Alkaissi and McFarlane, 2023). Despite our structured prompt, we needed to evaluate the reliability of ChatGPT’s output. Two coders manually reviewed a sample of explainable warnings generated for 100 randomly selected websites from our training dataset. Coders were provided with a screenshot of the website, its parsed source code, and its URL. They assessed: a) if GPT-3.5T’s features were present on the website, b) if the feature location was correct, c) if the feature description was accurate, and d) if any additional features were missed. The coders achieved a Cohen’s Kappa inter-reliability agreement of 0.78. After reviewing the websites, they resolved disagreements through discussion. In total, 401 feature points were identified across 100 websites. The targeted organization imitated by the phishing website was correctly identified in all cases. Coders found 383 ( 96%) features correctly identified, with the correct location for 339 ( 84%) of them. Among the 18 incorrect identifications were non-sensitive text fields (9), hallucinated two-factor authentication modules (5), and multiple brand texts (4).
7. Client-side extension
One of the primary motivations behind designing PhishLang was to create a detection framework that is both efficient and resource-intensive, allowing users to run the protection measures entirely on the client side without relying on online blocklists. Thus we build a client-side application that integrates a Chromium-based browser extension with a background process, that can block phishing websites without consulting any online blocklists. Once installed, the application automatically sets up our trained MobileBERT-based model as a local server, which the browser extension leverages to analyze all visited URLs. This setup is compatible with Microsoft Windows 10 and later versions, as well as Debian-based Linux distributions like Ubuntu (with MacOS support coming soon). The installation process configures the model, initiates the local service, and installs the extension for supported Chromium-based browsers. Operating entirely on CPU DRAM, the phishing detection processing is conducted entirely on the local machine.
This approach offers several advantages. Firstly, it ensures that the user’s browser history remains private and is not sent to any remote server. Prior studies have shown that users are often hesitant to share their browser history with third parties (Mangiò et al., 2020). Secondly, as discussed in Section 4.2, numerous studies, including our own, have demonstrated that online anti-phishing solutions frequently suffer from delays and inefficiencies when detecting both standard and evasive phishing threats. These delays can be particularly problematic when such solutions are integrated into anti-phishing browser extensions, significantly slowing down threat detection. By using our extension, users benefit from a robust detection model running directly on their own systems, thereby avoiding the delays associated with blocklists. Thirdly, as an open-source tool, our extension circumvents the need for substantial investment in server infrastructure, which would be necessary if detection were handled server-side. This setup not only limits scalability but also introduces risks of latency issues and downtimes. Running the extension entirely on the client side eliminates these concerns.
Also, recent research (Pourmohamad et al., 2024) has indicated that client-side implementations of in-browser anti-phishing measures, such as Google Safe Browsing, typically rely on simple heuristics and basic machine learning algorithms. These methods often lead to high false positive rates and low detection rates (14%) for threats when not using online blocklists. In contrast, our extension offers an open-source alternative that utilizes a more advanced and flexible model architecture. In Section X of the Appendix, we evaluate the performance and efficiency of our extension across various system configurations. Our findings indicate that the extension uses between 28MB and 247MB of RAM at peak load, which is a small fraction of the memory available in most modern systems. It can also run efficiently on low-end systems with only 1GB of RAM and a single-core processor while maintaining high precision (0.95) and recall (0.96) rates. To enhance performance, the application includes an optional built-in whitelist that users can choose to disable. Users also have the option to report URLs to our PhishLang database, though this feature is completely voluntary and does not prompt for user action. Additionally, users can add detected websites to a personal ’Safe List’ or report URLs as false positives to the PhishLang database if necessary. Figure 5 provides an illustration of the client-side extension in operation.
The Explainable Blocklisting feature is also available, though it is disabled by default. To enable this feature, users must provide their OpenAI key to access the explainable functionality, as detailed in Section 6. While this requires sharing the parsed representation of the website being visited, OpenAI allows users to prevent their provided prompts or generated responses used from being used for training or other purposes (OpenAI, 2024). We plan to conduct future research to offer this explainable feature on the client side as well, especially as more resource-efficient open-domain chatbots like Gemma become available (Google, 2024). However, these chatbots are not yet lightweight enough for practical local use on low to medium-end systems.
Figure 5. PhishLang client-side extension running in Brave Browser. Note that the website URL is not redacted since it is already inactive.
@@ 8. Conclusion
In this work, we introduced PhishLang, a novel phishing detection framework leveraging the MobileBERT architecture to identify phishing websites by comparing source codes with established phishing patterns. PhishLang outperforms various machine learning and deep learning models, reducing detection time and resources while showing strong resilience against adversarial attacks. In real-time, it identified several new phishing websites, including evasive attacks that were missed by popular anti-phishing tools. Reporting these threats improved detection rates, demonstrating PhishLang’s impact. We integrated PhishLang with GPT-3.5T for ”Explainable Blocklisting,” providing contextual explanations for phishing alerts, enhancing user understanding. We also implemented PhishLang as the first open-source client-side browser extension that does not require an external blocklist.